# Grouped Discrete Representation for Object-Centric Learning

### v2

@@ Abstract
Keywords:
Object-Centric Learning Variational Autoencoder Discrete Representation Channel Grouping.
@@ 1 Introduction
Under self or weak supervision, Object-Centric Learning (OCL) [10, 5] represents dense image or video pixels as sparse object feature vectors, known as slots. These slots can be used for set prediction while their corresponding attention maps for object discovery [20]. OCL is bio-plausible, as humans perceive visual scenes as objects for visual cognition, like understanding, reasoning, planning, and decision-making [2, 7, 22]. OCL is versatile, as object-level representations of images or videos fit to tasks involving different modalities [34, 31].
The training signal comes from reconstructing the input. Directly reconstructing input pixels [20, 10] struggles with complex-textured objects. Mixture-based OCL methods [16, 9] reconstruct more object-separable modalities, like optical flow and depth maps. Foundation-based methods [24, 35] use the input’s foundation model features as the target. Transformer-based [25, 27] and Diffusion-based methods [32, 14] reconstruct the input’s Variational Autoencoder (VAE) intermediate representation. With a limited number of shared template features, i.e., codes in a codebook, continuous-valued superpixels in VAE representations are discretized [12, 28]. This suppresses (super-)pixel noise and enhances object separability. Empirically, improved object separability in the reconstruction target offers OCL more effective training guidance.
@@
As illustrated in Fig. 1, consider a dataset characterized by two attribute groups: color (black, white) and shape (triangle, square, circle). An image in it contains four objects, each downsampled to a superpixel in the feature map. To select template features from a feature-level codebook, six scalar code indexes are needed, where digits 0-5 refer to black-triangle, black-circle, black-square, etc. Each code is reused with probability $\frac{1}{6}$. The feature map can thus be discretized as $[\begin{smallmatrix}0&4\\ 5&1\end{smallmatrix}]$. But if decomposed, superpixels can be discretized as combinations of template attributes from two attribute groups, i.e., $[\begin{smallmatrix}0,0&1,1\\ 1,2&0,1\end{smallmatrix}]$. The first and second numbers in these index tuples indicate whether superpixels’ attributes are the same or different, facilitating model convergence. These codes are reused with higher probabilities $\frac{1}{2}$ and $\frac{1}{3}$ respectively, benefiting model generalization.
@@ 2 Related Work
Object-Centric Learning (OCL). Mainstream OCL obtains supervision from reconstruction using slots aggregated by SlotAttention [20, 1] from the input’s dense superpixels. SLATE [25] and STEVE [27], which are Transformer-based, generate input tokens from slots via a Transformer decoder [29], guided by dVAE [12] discrete representations. SlotDiffusion [32] and LSD [14], which are Diffusion-based, recover input noise from slots via a Diffusion model [23], guided by VQ-VAE [28] discrete representations. DINOSAUR [24] and VideoSAUR [35], which are foundation model-based, reconstruct input features from slots via a spatial broadcast decoder [30], guided by well-pretrained features of the foundation model DINO [6, 21]. We focus on the VAE part of OCL.
Variational Autoencoder (VAE). Discrete representations of VAE have been shown to guide OCL better than direct input pixels as reconstruction targets. Transformer-based OCL methods [25, 27] utilize dVAE [12] to discretize encoder representations by selecting template features from a codebook via Gumbel sampling [13]. Diffusion-based OCL methods [14, 32] employ VQ-VAE [28] to achieve discretization by replacing features with their closest codebook codes. Similar to our idea, both [4] and [19] seek to decompose features into attributes, but their monolithic VAE representation is incompatible with OCL. Other VAE variants also offer techniques, like grouping [33], residual [3] and clustering [18], to enhance VAE representations. We borrow some for the OCL setting.
Channel Grouping. Splitting features along the channel dimension and transforming them separately is often used to diversify representations [17, 8, 11, 37, 36]. These solutions mainly perform grouping directly on feature maps [17] or on learnable parameters [36]. To the best of our knowledge, only one work has explored this idea in the OCL setting. SysBinder [26] groups slots along the channel dimension in the slot attention [20] to aggregate different attributes of objects, yielding better interpretability in object representation but limited performance gains. We group VAE intermediate representations along channels, yielding grouped discrete representations to guide OCL training better.
@@ 3 Proposed Method
We propose Grouped Discrete Representation (GDR), applicable to mainstream OCL methods, either Transformer-based [25, 27, 15, 35] or Diffusion-based [32, 14]. Simply modifying their VAE, our GDR improves them by providing reconstruction targets, or guidance, with better object separability.
@@ 3.1 Preliminary: Discrete Representation
Transformer-based architecture is depicted in Fig. 2 first row left. The input image or video frame $\bm{I}$ is encoded by a primary encoder and aggregated by SlotAttention [20] into slots $\bm{S}$ under queries $\bm{Q}$, with object (and background) segmentation masks $\bm{M}$ as byproducts. Meanwhile, pretrained VAE represents $\bm{I}$ as discrete representation $\bm{X}$ and the corresponding code indexes $\bm{X}_{\mathrm{i}}$. Subsequently, using a Transformer decoder, $\bm{S}$ is tasked with reconstructing $\bm{X}_{\mathrm{i}}$ as classification, guided by causal-masked $\bm{X}$. For videos, current slots $\bm{S}$ are transformed by a Transformer encoder block into queries for the next frame.
@@ 3.2 Naive Grouped Discrete Representation
[…]
@@
$\small l_{\mathrm{u}}=-\mathrm{entropy}(\mathbb{E}[D^{(1)}_{\mathrm{s}}])-\mathrm{entropy}(\mathbb{E}[D^{(2)}_{\mathrm{s}}])...-\mathrm{entropy}(\mathbb{E}[D^{(g)}_{\mathrm{s}}])$ (6)
@@
Remark. As illustrated in Fig. 1, by decomposing features into more reusable attributes, ideally any feature can be represented as a combination of these attributes, thus enhancing generalization. By indexing features with tuples rather than scalars, attribute-level similarities and differences can be captured for better object separability, thus benefiting convergence. Notably, when $g$=1, the above formulation except Eq. 6 reduces to the original non-grouped VAE.
@@ 3.3 Organizing Channel Grouping
where $\bm{Z}$ is in shape (height, width, $c$); $\mathrm{pinv}(\cdot)$ is pseudo-inverse; and matrix $\mathrm{pinv}(\bm{W})$ is in shape ($c$, expanded channel=$8c$). This facilitates channels belonging to the same attribute to be placed together by (i) enabling channel reordering and (ii) generating extra channels to mitigate mis-grouping.
@@
where $\bm{W}$ is the previously introduced learnable matrix in shape (expanded $c$, $c$).
@@ 3.4 Grouped vs Non-Grouped
Codebook parameters. Compared to the non-grouped, the number of parameters in our grouped codebook is significantly reduced to $\frac{agd}{a^{g}c}$=$\frac{ac}{a^{g}c}$=$\frac{1}{a^{g-1}}$. E.g., only $\frac{1}{64}$ when $a$=64, $g$=2, $c$=256 and $a^{g}$=4096. We increase $c$ to $8c$ and apply normalization plus linear to project it back to $c$, yielding $\frac{1}{1.6}$ the original number of codebook parameters – still 30% fewer.
Codebook computation. Non-grouped computation only involves code matching using inner product for each continuous feature: $c\times n\times 1$=$256\times 4096$=$2^{20}$. GDR computation involves two projections and code matching: $8c\times c\times 2+8c\times\sqrt[g]{n}$, which results in $2^{20}+2^{17}$ for $g$2 and $2^{20}+2^{14}$ for $g$4 – computation burden that is nearly identical to the original non-grouped case.
@@ 4.1 Experiment Overview
COCO #slots=7 class@top1$\uparrow$ bbox@R2$\uparrow$
@@
Models. We use both Transformer-based and Diffusion-based models as our GDR’s basis. The former includes SLATE [25] for image and STEVE [27] for video. The latter includes SlotDiffusion [32] and its temporal variant. Upon such basis, we compare GDR against SysBinder@$g$4 [26]. We also apply GDR to state-of-the-art models, SPOT [15] and VideoSAUR [35], which are also Transformer-based. Methods such as SA [20] and SAVi [16] are excluded due to their low performance or reliance on additional modalities.
Datasets. We evaluate those models on ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex, COCO [^3] https://cocodataset.org/#panoptic-2020 and VOC [^4] http://host.robots.ox.ac.uk/pascal/VOC for image OCL tasks, while MOVi-C/D/E [^5] https://github.com/google-research/kubric/tree/main/challenges/movi for video. We also use YTVIS [^6] https://youtube-vos.org/dataset/vis, YouTube video instance segmentation. These encompass both synthetic and real-world cases, featuring multiple objects and complex textures. Except for those two state-of-the-arts, the input size is unified to 128$\times$128 and other data processing follows the convention. Note that we use COCO panoptic instead of instance segmentation and the high-quality YTVIS [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset for strict evaluation.
Hyperparameters. The codebook size is $n$=$a^{g}$=4096 for all. GDR’s group number is set to GDR@$g$2, $g$4, $g$8 and $g$12. Correspondingly, the attribute group sizes are (64, 64), (8, 8, 8, 8), (2, 2, 2, 2, 4, 4, 4, 4) and (2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2), ensuring 4096 combinations. The number of slots is roughly set to the maximum/average object count plus one: 10+1 for ClevrTex, COCO, VOC and MOVi-C; 20+1 for MOVi-D; and 23+1 for MOVi-E. However, for those two state-of-the-arts, we strictly follow their official experiment settings.
@@ 4.2 Performance
Object discovery. We use common object discovery metrics: Adjusted Rand Index (ARI) [^8] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARI foreground (ARIfg), mean Best Overlap (mBO) [^9] https://ieeexplore.ieee.org/document/7423791 and mean Intersection-over-Union (mIoU) [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html. As shown in Fig. 5, GDR significantly enhances accuracy across both synthetic and real-world images and videos. With the naive CNN [16] for primary encoding, it boosts both Transformer-based methods and Diffusion-based methods. GDR always outperforms the competitor SysBinder by a large margin. We further evaluate GDR’s effectiveness with vision foundation model DINO1-B/8 [6] for strong primary encoding. As shown in Fig. 6 and 3, on both SLATE and SlotDiffusion, GDR improves accuracy across all metrics in most cases.
Applying to state-of-the-art. Following the original settings, we apply GDR to SPOT [15] and VideoSAUR [35], by replacing SPOT’s VAE with GDR and by replacing VideoSAUR’s reconstruction target (continuous DINO features) with GDR discretized DINO features, respectively. As shown in Table 1 upper, GDR is still able to boost state-of-the-art methods’ performance further.
Set prediction. Following [24], we employ OCL to represent dataset COCO as slots, and use a small MLP to predict the object class and bounding box corresponding to each slot, with metrics of top-1 accuracy and the R2 score respectively. As shown in Table 1 lower, our GDR improves SLATE in set prediction, demonstrating is superior quality in object representation.
@@
Figure 8: For GDR@$g$2, one attribute group roughly learns colors, while the other roughly learns textures. The original image is at the center. The left and right are images decoded from the modified VAE discrete representation.
@@ 4.3 Interpretability
Decomposition from features to attributes. Although without explicit supervision models can hardly learn concepts [26] as human-readable as in Fig. 1, we can still analyze GDR’s decomposition from features to attributes as follows. Given GDR discrete representation $\bm{X}$, we replace the attributes of objects’ superpixels with arbitrary attribute codes then decode them into images to observe the changes. As shown in Fig. 8, under $g$2 setting, modifying one attribute group roughly alters the colors, whereas modifying the other destroys the textures. This suggests that the first group learns colors while the second learns textures.
@@ 4.4 Ablation
Number of groups, formulated in Eq. 1-6: $g$=2, 4, 8 or 12. As shown in Fig. 5 and 6, the optimal $g$ depends on the specific dataset. However, $g$12 and $g$8 tend to result in suboptimal performance while $g$4 consistently leads to guaranteed performance gains over the basis methods.
[…]
@@
utiliz. loss invertible project residual connection final normaliz. ARI+ARIfg 89.47 84.78 81.52 $W$pinv 32.25 88.84 89.16
Table 2: Effects of expansion rate, utilization loss, invertible projection (and replacing $W$ pinv with specified weights), residual connection in training and final normalization. Experimented SLATE+GDR@$g$4 on ClevrTex.

### v3

@@ 1 Introduction
Under self or weak supervision, Object-Centric Learning (OCL) [greff2019iodine, burgess2019monet] represents dense image or video pixels as sparse object feature vectors, known as slots. These slots can be used for set prediction while their corresponding attention maps for object discovery [locatello2020slotattent]. OCL is bio-plausible, as humans perceive visual scenes as objects for visual cognition, like understanding, reasoning, planning, and decision-making [bar2004visual, cavanagh2011visual, palmeri2004visual]. OCL is versatile, as object-level representations of images or videos fit to tasks involving different modalities [yi2019clevrer, wu2022slotformer].
The training signal comes from reconstructing the input. Directly reconstructing input pixels [locatello2020slotattent, greff2019iodine] struggles with complex-textured objects. Mixture-based OCL methods [kipf2021savi, elsayed2022savipp] reconstruct more object-separable modalities, like optical flow and depth maps. Foundation-based methods [seitzer2023dinosaur, zadaianchuk2024videosaur] use the input’s foundation model features as the target. Transformer-based [singh2021slate, singh2022steve] and Diffusion-based methods [wu2023slotdiffuz, jiang2023lsd] reconstruct the input’s Variational Autoencoder (VAE) intermediate representation. With a limited number of shared template features, i.e., codes in a codebook, continuous-valued superpixels in VAE representations are discretized [im2017dvae, van2017vqvae]. This suppresses (super-)pixel noise and enhances object separability. Empirically, improved object separability in the reconstruction target offers OCL more effective training guidance.
@@
As illustrated in Fig. 1, consider a dataset characterized by two attribute groups: color (black, white) and shape (triangle, square, circle). An image in it contains four objects, each downsampled to a superpixel in the feature map. To select template features from a feature-level codebook, six scalar code indexes are needed, where digits 0-5 refer to black-triangle, black-circle, black-square, etc. Each code is reused with probability $\frac{1}{6}$. The feature map can thus be discretized as $\begin{smallmatrix}0&4\\ 5&1\end{smallmatrix}]$. But if decomposed, superpixels can be discretized as combinations of template attributes from two attribute groups, i.e., $\begin{smallmatrix}0,0&1,1\\ 1,2&0,1\end{smallmatrix}]$. The first and second numbers in these index tuples indicate whether superpixels’ attributes are the same or different, facilitating model convergence. These codes are reused with higher probabilities $\frac{1}{2}$ and $\frac{1}{3}$ respectively, benefiting model generalization.
@@ 2 Related Work
Object-Centric Learning (OCL). Mainstream OCL obtains supervision from reconstruction using slots aggregated by SlotAttention [locatello2020slotattent, bahdanau2014qkvattent] from the input’s dense superpixels. SLATE [singh2021slate] and STEVE [singh2022steve], which are Transformer-based, generate input tokens from slots via a Transformer decoder [vaswani2017transformer], guided by dVAE [im2017dvae] discrete representations. SlotDiffusion [wu2023slotdiffuz] and LSD [jiang2023lsd], which are Diffusion-based, recover input noise from slots via a Diffusion model [rombach2022ldm], guided by VQ-VAE [van2017vqvae] discrete representations. DINOSAUR [seitzer2023dinosaur] and VideoSAUR [zadaianchuk2024videosaur], which are foundation model-based, reconstruct input features from slots via a spatial broadcast decoder [watters2019sbd], guided by well-pretrained features of the foundation model DINO [caron2021dino1, oquab2023dino2]. We focus on the VAE part of OCL.
Variational Autoencoder (VAE). Discrete representations of VAE have been shown to guide OCL better than direct input pixels as reconstruction targets. Transformer-based OCL methods [singh2021slate, singh2022steve] utilize dVAE [im2017dvae] to discretize encoder representations by selecting template features from a codebook via Gumbel sampling [jang2016gumbelsoftmax]. Diffusion-based OCL methods [jiang2023lsd, wu2023slotdiffuz] employ VQ-VAE [van2017vqvae] to achieve discretization by replacing features with their closest codebook codes. Similar to our idea, both [bouchacourt2018mlvae] and [liu2023cdisentangle] seek to decompose features into attributes, but their monolithic VAE representation is incompatible with OCL. Other VAE variants also offer techniques, like grouping [yang2023hificodec], residual [barnes1996residualvq] and clustering [lim2020clustervae], to enhance VAE representations. We borrow some for the OCL setting.
Channel Grouping. Splitting features along the channel dimension and transforming them separately is often used to diversify representations [krizhevsky2012alexnet, chen2019octconv, huang2018condensenet, zhao2021lhc, zhao2022coc]. These solutions mainly perform grouping directly on feature maps [krizhevsky2012alexnet] or on learnable parameters [zhao2022coc]. To the best of our knowledge, only one work has explored this idea in the OCL setting. SysBinder [singh2022sysbind] groups slots along the channel dimension in the slot attention [locatello2020slotattent] to aggregate different attributes of objects, yielding better interpretability in object representation but limited performance gains. We group VAE intermediate representations along channels, yielding grouped discrete representations to guide OCL training better.
@@ 3 Proposed Method
We propose Grouped Discrete Representation (GDR), applicable to mainstream OCL methods, either Transformer-based [singh2021slate, singh2022steve, kakogeorgiou2024spot, zadaianchuk2024videosaur] or Diffusion-based [wu2023slotdiffuz, jiang2023lsd]. Simply modifying their VAE, our GDR improves them by providing reconstruction targets, or guidance, with better object separability.
@@ 3.1 Preliminary: Discrete Representation
Transformer-based architecture is depicted in Fig. 2 first row left. The input image or video frame $\bm{I}$ is encoded by a primary encoder and aggregated by SlotAttention [locatello2020slotattent] into slots $\bm{S}$ under queries $\bm{Q}$, with object (and background) segmentation masks $\bm{M}$ as byproducts. Meanwhile, pretrained VAE represents $\bm{I}$ as discrete representation $\bm{X}$ and the corresponding code indexes $\bm{X}_{\mathrm{i}}$. Subsequently, using a Transformer decoder, $\bm{S}$ is tasked with reconstructing $\bm{X}_{\mathrm{i}}$ as classification, guided by causal-masked $\bm{X}$. For videos, current slots $\bm{S}$ are transformed by a Transformer encoder block into queries for the next frame.
@@ 3.2 Naive Grouped Discrete Representation
where $\bm{Z}^{(1)}$$\bm{Z}^{(g)}$ are channel groupings of $\bm{Z}$; $\bm{G}^{(1)}$$\bm{G}^{(g)}$ are Gumbel noises; $\circ$ is channel concatenation; $\mathrm{l2}(\cdot,\cdot)$ denotes L2 distances between every vector pair in its two arguments; $\bm{D}_{\mathrm{s}}$ is soft Gumbel sampling of distances $\bm{D}$ between continuous representations and codes; $\mathrm{argmin}(\cdot)$ is along the code dimension. For our grouped VAE, multiple code indexes are selected from the attribute groups, forming “tuple indexes”. In contrast, the non-grouped VAE selects only one index from a feature-level codebook, forming “scalar indexes”.
@@
[…]
@@
Remark. As illustrated in Fig. 1, by decomposing features into more reusable attributes, ideally any feature can be represented as a combination of these attributes, thus enhancing generalization. By indexing features with tuples rather than scalars, attribute-level similarities and differences can be captured for better object separability, thus benefiting convergence. Notably, when gg=1, the above formulation except Eq. 6 reduces to the original non-grouped VAE.
@@ 3.3 Organizing Channel Grouping
where $\bm{Z}$ is in shape (height, width, channel=$c$); $\mathrm{pinv}(\cdot)$ is pseudo-inverse; and matrix $\mathrm{pinv}(\bm{W})$ is in shape (channel=$c$, expanded channel=$8c$). This facilitates channels belonging to the same attribute to be placed together by (i) enabling channel reordering and (ii) generating extra channels to mitigate mis-grouping.
@@
where $\bm{W}$ is the previously introduced learnable matrix in shape (expanded channel=8cc, channel=$c$).
@@ 3.4 Grouped vs Non-Grouped
Codebook parameters. Compared to the non-grouped, the number of parameters in our grouped codebook is significantly reduced to $\frac{agd}{a^{g}c}$$\frac{ac}{a^{g}c}$$\frac{1}{a^{g-1}}$. E.g., only $\frac{1}{64}$ when aa=64, gg=2, cc=256 and $a^{g}$=4096. We increase $c$ to $8c$ and apply normalization plus linear to project it back to $c$, yielding $\frac{1}{1.6}$ the original number of codebook parameters – still 30% fewer.
Codebook computation. Non-grouped computation only involves code matching using inner product for each continuous feature: $c\times n\times 1$$56\times 4096=22$$^{20}$. GDR computation involves two projections and code matching: $8c\times c\times 2+8c\times\sqrt[g]{n}$, which results in $2^{20}+2^{17}$ for gg2 and $2^{20}+2^{14}$ for gg4 – computation burden that is nearly identical to the original non-grouped case.
@@ 4.1 Experiment Overview
[…]
@@
Models. We use both Transformer-based and Diffusion-based models as our GDR’s basis. The former includes SLATE [singh2021slate] for image and STEVE [singh2022steve] for video. The latter includes SlotDiffusion [wu2023slotdiffuz] and its temporal variant. Upon such basis, we compare GDR against SysBinder@gg4 [singh2022sysbind]. We also apply GDR to state-of-the-art models, SPOT [kakogeorgiou2024spot] and VideoSAUR [zadaianchuk2024videosaur], which are also Transformer-based. Methods such as SA [locatello2020slotattent] and SAVi [kipf2021savi] are excluded due to their low performance or reliance on additional modalities.
Datasets. We evaluate those models on ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex, COCO [^3] https://cocodataset.org/#panoptic-2020 and VOC [^4] http://host.robots.ox.ac.uk/pascal/VOC for image OCL tasks, while MOVi-C/D/E [^5] https://github.com/google-research/kubric/tree/main/challenges/movi for video. We also use YTVIS [^6] https://youtube-vos.org/dataset/vis, YouTube video instance segmentation. These encompass both synthetic and real-world cases, featuring multiple objects and complex textures. Except for those two state-of-the-arts, the input size is unified to $\times128$ and other data processing follows the convention. Note that we use COCO panoptic instead of instance segmentation and the high-quality YTVIS [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset for strict evaluation.
[…]
@@ 4.2 Performance
Object discovery. We use common object discovery metrics: Adjusted Rand Index (ARI) [^8] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARI foreground (ARIfg), mean Best Overlap (mBO) [^9] https://ieeexplore.ieee.org/document/7423791 and mean Intersection-over-Union (mIoU) [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html. As shown in Fig. 5, GDR significantly enhances accuracy across both synthetic and real-world images and videos. With the naive CNN [kipf2021savi] for primary encoding, it boosts both Transformer-based methods and Diffusion-based methods. GDR always outperforms the competitor SysBinder by a large margin. We further evaluate GDR’s effectiveness with vision foundation model DINO1-B/8 [caron2021dino1] for strong primary encoding. As shown in Fig. 6 and 3, on both SLATE and SlotDiffusion, GDR improves accuracy across all metrics in most cases.
Applying to state-of-the-art. Following the original settings, we apply GDR to SPOT [kakogeorgiou2024spot] and VideoSAUR [zadaianchuk2024videosaur], by replacing SPOT’s VAE with GDR and by replacing VideoSAUR’s reconstruction target (continuous DINO features) with GDR discretized DINO features, respectively. As shown in Table 1 upper, GDR is still able to boost state-of-the-art methods’ performance further.
Set prediction. Following [seitzer2023dinosaur], we employ OCL to represent dataset COCO as slots, and use a small MLP to predict the object class and bounding box corresponding to each slot, with metrics of top-1 accuracy and the R2 score respectively. As shown in Table 1 lower, our GDR improves SLATE in set prediction, demonstrating is superior quality in object representation.
@@
Figure 8: For GDR@gg2, one attribute group roughly learns colors, while the other roughly learns textures. The original image is at the center. The left and right are images decoded from the modified VAE discrete representation.
@@ 4.3 Interpretability
Decomposition from features to attributes. Although without explicit supervision models can hardly learn concepts [singh2022sysbind] as human-readable as in Fig. 1, we can still analyze GDR’s decomposition from features to attributes as follows. Given GDR discrete representation $\bm{X}$, we replace the attributes of objects’ superpixels with arbitrary attribute codes then decode them into images to observe the changes. As shown in Fig. 8, under gg2 setting, modifying one attribute group roughly alters the colors, whereas modifying the other destroys the textures. This suggests that the first group learns colors while the second learns textures.
@@ 4.4 Ablation
Number of groups, formulated in Eq. 1-6: gg=2, 4, 8 or 12. As shown in Fig. 5 and 6, the optimal $g$ depends on the specific dataset. However, gg12 and gg8 tend to result in suboptimal performance while gg4 consistently leads to guaranteed performance gains over the basis methods.
Channel expansion rate, formulated in Eq. 7: $c$, 2cc, 4cc or 8cc. Although $8c$ generally performs best, the expansion rate has a nearly saturated impact on GDR’s performance. This suggests that our channel organizing mechanism is effective, reducing the necessity for a higher channel expansion rate.
@@
utiliz. loss invertible project residual connection final normaliz. ARI+ARIfg 89.47 84.78 81.52 WWpinv 32.25 88.84 89.16
Table 2: Effects of expansion rate, utilization loss, invertible projection (and replacing $W$ pinv with specified weights), residual connection in training and final normalization. Experimented SLATE+GDR@gg4 on ClevrTex.